\begin{thebibliography}{99}

\bibitem{Arnold}
D. Arnold, \emph{Finite-Rank Torsion Free Abelian Groups and Rings}, Lecture Notes in Mathematics, 931, Springer-Verlag, (1982).

\bibitem{DikranjanLewisLothMader}
D. Dikranjan, W. Lewis, P. Loth, and A. Mader, \emph{A distinguished subgroup of compact abelian groups}, Axioms 11(5), 200, (2022).

\bibitem{Fuchs}
L. Fuchs, \emph{Abelian Groups}, Springer Monographs in Mathematics, Springer, Switzerland, 2015.

\bibitem{HHR}
W. Herfort, K. Hofmann, F. Russo, \emph{Periodic Locally Compact Groups}, de Gruyter, 2019.

\bibitem{HofmannMorris}
K. H. Hofmann and S. A. Morris, \emph{The Structure of Compact Groups, a Primer for the Student---A Handbook for the Expert}, 4th ed.; De Gruyter: Berlin, Germany, 2020.

\bibitem{Lewis2019}
W. Lewis, \emph{Structure of Finite-Dimensional Protori}, Axioms, 8(3), 93, 2019.

\bibitem{Moskowitz}
M. Moskowitz, \emph{Homological Algebra in Locally Compact Abelian Groups}, Transactions of the American Mathematical Society, 127(3), (1967), p.361--404.

\bibitem{RZ}
L. Ribes and P. Zalesskii, \emph{Profinite Groups}, 2nd ed., Ergebnisse der Mathematik und ihrer Grenzgebiete, Volume 40, Springer, Berlin/Heidelberg, 2010.

\bibitem{Zassenhaus}
H. Zassenhaus, \emph{Orders as Endomorphism Rings of Modules of the Same Rank}, Journal of the London Mathematical Society, 42, (1967), p. 180--182.

\end{thebibliography}

\documentclass[10pt,letterpaper]{amsart}

\usepackage[letterpaper, margin=1in]{geometry}
\usepackage{amssymb,amsthm,amsmath}
\usepackage{mathtools}
\usepackage[utf8]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{tikz}
\usepackage{enumitem}
\usepackage{changepage}
\usepackage{tikz-cd}
\usepackage{textcomp}
\DeclareUnicodeCharacter{02BB}{\textquoteleft}
\usepackage{graphicx}
\usepackage[final]{microtype}
\usepackage[section]{placeins}
\usepackage{etoolbox}
\usepackage[numbers]{natbib}
\usepackage[pdftex,unicode]{hyperref}
\usepackage{lmodern}

\usetikzlibrary{arrows.meta, positioning}

\theoremstyle{plain}
\newtheorem{thm}[equation]{Theorem}
\newtheorem{lem}[equation]{Lemma}

\theoremstyle{definition}

\newtheorem{example}[equation]{Example}

\theoremstyle{remark}
\newtheorem{rem}[equation]{Remark}

\newcommand{\Z}{\mathbb{Z}}
\newcommand{\N}{\mathbb{N}}

\newcommand{\A}{\mathbb{A}}
\newcommand{\az}{\A_{\z}}
\newcommand{\q}{\mathbb{Q}}
\newcommand{\z}{\mathbb{Z}}
\newcommand{\ev}{\operatorname{ev}}

\newcommand{\F}{\mathbb{F}}

\newcommand{\p}{\mathbb{P}}

\newcommand{\lowfrown}[1]{\overset{\raisebox{-0.35ex}{$\scriptstyle\frown$}}{#1}}

\newcommand{\bt}{\mathbf{t}}

\newcommand{\zhat}{\widehat{\Z}}
\newcommand{\nala}{\mathcal{L}_{\rm n}}
\newcommand{\arla}{\mathcal{L}_{\rm a}}
\newcommand{\adla}{\mathcal{L}}

\newcommand{\qhat}{\widehat{\q}}

\newcommand{\dov}{\mathrm{d}}

\newcommand{\ra}{\rightarrow}
\newcommand{\deff}{:=}

\newcommand{\OO}{\mathcal{O}}

\newcommand{\R}{\mathbb{R}}

\renewcommand{\emph}[1]{{\it #1}}

\DeclareMathOperator{\id}{id}

\DeclareMathOperator{\End}{End}

\DeclareMathOperator{\Hom}{Hom}

\DeclareMathOperator{\G}{G}

\DeclareMathOperator{\lcm}{lcm}
\DeclareMathOperator{\td}{td}

\DeclareMathOperator{\rk}{rk}
\DeclareMathOperator{\dima}{dim_a}
\DeclareMathOperator{\dimn}{dim_n}

\title[Finite-Dimensional Protori Are Adelic Tori]{Finite-Dimensional Protori Are Adelic Tori}
\author{Wayne Lewis}
\address{University of Hawaiʻi, Honolulu Community College}
\email{waynel@math.hawaii.edu}
\date{\today}

\subjclass[2020]{Primary 22B05; Secondary 22C05, 20K15, 22E40, 12F05}

\begin{document}
	
\begin{abstract}
We show that the category of finite-dimensional compact connected abelian groups is naturally realised as the category of adelic tori $\frac{(\A^n/C)}{\Lambda}$ for the adeles $\A$, a profinite group $C$, and a discrete cocompact $\Lambda\cong\q^n$.
A Lie theory is developed yielding a proper short exact sequence $0\ra \q^n \ra \mathcal{L}(G)\xrightarrow{\exp}\G\ra 0$ for each adelic torus $G$ where $\exp$ denotes the adelic exponential and $\mathcal{L}(G)$ the adelic Lie algebra of $G$.
The perfect symmetry of the adelic structure theory of the category $\mathbf{AT}$ of adelic tori under continuous quasi-homomorphism enables us to establish that the nonarchimedean dimension of an arbitrary number field $K$ with $[K\colon \q]=n$ is $1$, while the archimedean dimension is $n$.
\end{abstract}
\maketitle

From its inception, the structure theory of compact connected abelian groups (also known as pro-tori \cite[Definition 9.30]{HofmannMorris}, herein contracted to protori) has been biased toward archimedean invariants, treating totally disconnected components as pathological or secondary.

By recasting compact connected abelian groups explicitly as adelic tori, we recover a perfect structural symmetry between their nonarchimedean and archimedean geometries. 
This adelic perspective not only clarifies the topological boundaries of these spaces but naturally connects their continuous geometry to the discrete arithmetic of number fields and the Krull-Schmidt category of discrete torsion-free abelian groups under quasi-homomorphism.

The purpose of this paper is to establish the foundational adelic machinery for these spaces; we proceed directly to the main results.
Let $\mathbf{LCA}$ denote the category of locally compact abelian groups under continuous homomorphisms, and let $H^\vee$ denote the Pontryagin dual of $H\in\mathbf{LCA}$. 
Define the {\it torsion-free rank} of a discrete finite-rank torsion-free abelian group $A$ by $\rk A\deff \dim_{\q}(\q\otimes_{\z}A)$; all tensor products herein are over $\z$. 
Define the {\it archimedean dimension} of a protorus $G$ by $\dima G\deff \rk(G^\vee)$ \cite[Definition 8.23]{HofmannMorris}; a solenoid is a protorus of archimedean dimension $1$.
We use the term {\it adelic torus} as a topological synonym for a finite-dimensional compact connected abelian group.

{\it Unless otherwise specified, all groups are abelian, and the symbol $\cong$ denotes topological isomorphism when it appears between topological groups.}

Define groups $A$ and $B$ to be quasi-isomorphic if there is $n\in\N$, $f\colon A\ra B$, and $g\colon B\ra A$ such that $fg=n\cdot \id_B$ and $gf=n\cdot \id_A$. 
A quasi-homomorphism between $A$ and $B$ is defined to be an element of $\q\otimes\Hom(A,B)$. Finite-rank torsion-free groups under quasi-homomorphism form an additive Krull-Schmidt category \cite[Theorem 7.4, Example 7.6]{Arnold}. 
Restriction of Pontryagin duality to discrete finite-rank torsion-free groups under quasi-homomorphism yields a contravariant equivalence with the category of finite-dimensional compact connected groups under
continuous quasi-homomorphism.

Let $G$ and $H$ be protori of archimedean dimension $n$. 
Define an {\it isogeny} from $G$ to $H$ to be a surjective continuous homomorphism $G\ra H$ with finite kernel. 
We say that $G$ and $H$ are {\it isogenous}, and write $G\simeq H$, if there exists an isogeny $G\ra H$. 
One shows that being isogenous is an equivalence relation and that there exists a continuous quasi-isomorphism $G\ra H$ if and only if there exists an isogeny $G\ra H$.
A topological group $K$ is defined to be an $n$-dimensional real torus, $0\leq n\in\Z$, if and only if $K\cong (\frac{\R}{\z})^n$ with the usual Euclidean topology on $\R$, subspace topology on $\Z$, quotient topology on $\frac{\R}{\z}$, and product topology on $(\frac{\R}{\z})^n$. 
A delta-subgroup $C$ of a protorus $G$ is a profinite subgroup such that $\frac{G}{C}$ is an $n$-dimensional real torus.

See \cite[p.6 \& Corollary 4]{Lewis2019} for the definition of {\it nonarchimedean dimension of $G$}, denoted $\dimn G$; the key property is that $\dimn G=\dimn C$ for a delta-subgroup $C$ of $G$.
Define the {\it adelic dimension of $G$} to be the ordered pair $(\dimn(G),\dima(G))$.

\begin{thm}[Dimension Extremes]\label{dimension-extremes}
Let $G$ be a protorus with adelic dimension $(m,n)=(\dimn G,\dima G)$. 
Then $m=0$ if and only if $G$ is an $n$-dimensional real torus, and $m=n$ if and only if $G\simeq \prod_{i=1}^n S_i$ is a product of solenoids, $1\leq i\leq n$, with no real torus factors.
\end{thm}
\begin{proof}
$G/\{0\}\cong G\cong (\R/\Z)^n \Leftrightarrow \{0\}$ is a delta-subgroup of $G\Leftrightarrow m=0$ \cite[Corollary 4]{Lewis2019}.

$m=n \Rightarrow$ the Pontryagin dual $G^\vee$ of $G$ is quasi-isomorphic to a rank-$n$ completely decomposable group with no free summands \cite[Corollary 10]{Lewis2019} $\Rightarrow G$ is isogenous to a product of $n$ solenoids with no torus factors $\Rightarrow G$ is a product of $n$ solenoids with no torus factors in the category of finite-dimensional compact connected groups under continuous quasi-homomorphism.
Conversely, additivity of $\dim_n$ over products of finite-dimensional protori gives $m=n$.
\end{proof}

A {periodic $\mathbf{LCA}$ group $L$} is an $\mathbf{LCA}$ group equal to the directed union of its compact open subgroups; let $L_p$ denote the $p$-Sylow component of $L$ \cite[Definition 2.2]{HHR}. 
A periodic $\mathbf{LCA}$ group $L$ with compact open subgroup $C$ is topologically isomorphic to the restricted product $\prod'_{p\in\p}(L_p,C_p)$ \cite[Theorem 3.3]{HHR}.

Let $\zhat$ denote the additive topological group of the ring of profinite integers, defined as the inverse limit of the system $\{\frac{\z}{n\z}\colon 1\le n\in\Z\}$ directed by divisibility; then $\zhat\cong \prod_{p\in\p}\Z_p$ where $\Z_p$ denotes the ring of $p$-adic integers \cite[Example 2.3.11]{RZ}. 
Let $\qhat = \prod'_{p\in\p}(\q_p,\z_p)$ denote the additive topological group of the locally compact ring of finite adeles; then $\qhat\cong \q\otimes\zhat$ \cite[Theorem 3.51]{HHR}. 
Let $\A=\qhat\times\R$ denote the additive topological group of the locally compact ring of adeles; then $\A/\q\cong \q^\vee$ is the universal solenoid: it is the only torsion-free solenoid and every solenoid is topologically isomorphic to a quotient of $\A/\q$, whence a quotient of $\A$ (by Pontryagin duality since every rank-one discrete torsion-free abelian group is isomorphic to an additive subgroup of $\q$).

\begin{thm}[Finite-Dimensional Compact Connected Groups Are Adelic Tori]\label{fdccg-adelic-tori}
If $G$ is a protorus with $\dima G=n$, then $G\cong (\A^n/C)/\Lambda$ for a profinite group $C$ and a discrete cocompact $\Lambda\cong \q^n$.
\end{thm}
\begin{proof}
Let $G$ be an $n$-dimensional compact connected group. 
By \cite[Theorem 1.3, $Y=\q^n$]{Lewis2019}, $G\cong (D\times\R^n)/\q^n$ where $D$ is a divisible periodic LCA group of nonarchimedean dimension $n$, and $\q^n$ is embedded diagonally as a discrete cocompact subgroup. 
By the structure theorem for periodic $\mathbf{LCA}$ groups \cite[Theorem 3.3]{HHR}, there is a (necessarily torsion-free) profinite subgroup $C\subseteq \qhat^n$ such that $D\cong \qhat^n/C$. 
Hence, $G\cong \frac{(\qhat^n/C)\times \R^n}{\q^n}\cong \frac{(\qhat^n\times \R^n)/C}{\q^n}
\cong \frac{\A^n/C}{\q^n}$ where $C$ is viewed as a subgroup of $\A^n=\qhat^n\times\R^n$. 
Thus, an $n$-dimensional compact connected group is topologically isomorphic to $(\A^n/C)/\q^n$ for $n,C$ as in Theorem \ref{fdccg-adelic-tori}.
\end{proof}

Define a {\it discrete torus} to be a discrete finite-rank torsion-free abelian group. 
Define the elements of $\q\otimes\Hom(A,B)$ for discrete tori $A$ and $B$ to be the quasi-homomorphisms $A\ra B$. 
For discrete tori $A$ and $B$ it is common practice in abelian group theory to abbreviate $\q\otimes A$ as $\q A$ and we make the identification $\q\otimes\Hom(A,B)=\q\Hom(A,B)$. 
Let $\mathbf{DT}$ denote the category of finite-rank torsion-free discrete tori under quasi-homomorphism.

Every abelian group $A$ decomposes as $A=\dov(A)\oplus M$ for the unique maximal divisible subgroup $\dov(A)$ and some complementary summand $M$ \cite[Theorem 2.5]{Fuchs}. 
Fix $B\in \mathbf{DT}$. 
Then $B$ is {\it locally free} if the localisation $B_{(p)}\cong\z_{(p)}\otimes_\z B$ (\cite[p.4]{Arnold}) is isomorphic to $\Z^{\rk B}_{(p)}$ for $p\in\p$; equivalently, $\dim_{\,\F_p}(\frac{B}{pB})=\rk B$ for $p\in\p$ \cite[p.16]{Arnold}. 
And $B$ is {\it quotient divisible} if there is $\Z^{\rk B}\cong F\subseteq B$ with ${\rm d}(\frac{B}{F})=\frac{B}{F}$ \cite[p.18]{Arnold}.

The set of {\it supernatural numbers $\mathcal S$} is defined to be the set of formal products $\prod_{p\in\p}p^{k_p}$, $0\leq k_p\leq \infty$. 
Define $\sim$ on $\mathcal S$ by $\bt\sim \bt'$ if and only if $k_p=\infty\Leftrightarrow k'_p=\infty$ and $k_p=k'_p$ for all but finitely many $p\in\p$; then $\sim$ is an equivalence relation on $\mathcal S$; the elements of $\frac{\mathcal S}{\sim}$, called (Baer) {\it types}, are partially ordered by $\bt\leq \bt'$ if and only if $k_p=\infty\Rightarrow k'_p=\infty$ and $k_p\leq k'_p$ for all but finitely many $p\in\p$; and one shows $\mathcal B\deff \frac{\mathcal S}{\sim}$ is a distributive lattice, called the Baer lattice, where join $\vee$ is induced by $\lcm$ and meet $\wedge$ is induced by $\gcd$.

Fix a rank-$1$ torsion free group $A$; fix a rational group $B$ (i.e., a group $B$ with $\Z\subseteq B\subseteq\q$) with $A\cong B$. 
Define the type of a nonzero element $a\in A$ to be $\bt_A(a)\deff \frac{\prod_{p\in\p}p^{h_p}}{\sim}\in\mathcal B$ where $h_p\deff \sup\{\ell\in\N\cup\{0\}\colon a/p^\ell\in A\}$, $p\in\p$. 
The (Baer) {\it type} of $A$ given by $\bt(A)\deff \bt_B(1)$ is well-defined. 
A torsion-free group $E$ is $\bt(B)$-homogeneous if $\bt_E(a)=\bt(B)$ for all $0\neq a\in E$.

A topological group $G$ is an {\it adelic torus} if $G\cong (\A^n/C)/\q^n$ where $\q^n$ denotes its diagonally embedded image in $\A^n/C$ for the (additive topological group of the) ring of adeles $\A$, some $0\leq n\in\Z$, and $C$ a torsion-free profinite subgroup of $\A^n$. 
Let $\mathbf{AT}$ denote the category of adelic tori under continuous quasi-homomorphism; then Pontryagin duality restricts to a contravariant equivalence between the category $\mathbf{DT}$ and the category $\mathbf{AT}$. 
Set $\mathcal{D}(G)\deff\{C\colon C \text{ a delta-subgroup of } G\}$ for $G\in\mathbf{AT}$.

\begin{thm}[Spectrum of Resolutions]\label{spectrum-resolutions}
Let $G\in\mathbf{AT}$, $\dim_a G=n$, and suppose $G\cong \frac{C'\times\R^n}{\z^n}$ for some delta-subgroup $C'\in \mathcal{D}(G)$. 
If $\z^n\subseteq B\subseteq \q^n$, then $G\cong \frac{L_B\times\R^n}{B}$, where $L_B=C'\oplus_{\mathbb Z^n}B$ is the {\rm (}periodic $\mathbf{LCA}${\rm )} pushout of $C'$ along $\mathbb Z^n\hookrightarrow B$ and $\Z^n\to C'$ is the projection of the diagonal embedding $\z^n\hookrightarrow C'\times\R^n$ onto its first coordinate.
\end{thm}
\begin{proof}
The existence of such a resolution associated with $B$ follows from the Structure Theorem for Finite-Dimensional Protori
\cite[Theorem 1.3, $Y=B$]{Lewis2019}. 
The following direct argument additionally identifies its periodic factor $L_B$ canonically, relative to the fixed resolution, as the stated pushout.
	
Let $\alpha\colon\z^n\ra C'$ be the projection onto the first coordinate of the diagonal embedding defining the given resolution, so that $G\cong	\frac{C'\times\R^n}{\{(\alpha(z),z)\colon z\in\z^n\}}$.
Give $B$ the discrete topology and realize the pushout as $L_B=\frac{C'\oplus B}{\{(\alpha(z),-z)\colon z\in\z^n\}}$, with its quotient topology. 
The subgroup $\{(\alpha(z),-z)\colon z\in\z^n\}$ is closed in
$C'\oplus B$, since convergence in the discrete second coordinate forces $z$ to be eventually constant; hence, $L_B$ is an $\mathbf{LCA}$ group.
Write $\iota_B(b)=[0,b]\in L_B$, and embed $B$ diagonally in $L_B\times\R^n$ by $b\mapsto(\iota_B(b),b)$.
	
Define $\Phi_B\colon L_B\times\R^n\ra G$ by $\Phi_B([c,b],x)=[(c,x-b)]$.
The defining relation $[c+\alpha(z),b-z]=[c,b]$	shows that $\Phi_B$ is well-defined. 
$\Phi_B$ is continuous and surjective and $\Phi_B(\iota_B(b),b)=0$ for every $b\in B$.
If $\Phi_B([c,b],x)=0$, then $(c,x-b)=(\alpha(z),z)$ for some $z\in\z^n$; hence, $x=b+z\in B$ and $[c,b]=[\alpha(z),b]=[0,b+z]=\iota_B(x)$; therefore, $\ker\Phi_B=\{(\iota_B(b),b)\colon b\in B\}$.
Thus, $\Phi_B$ induces a continuous isomorphism $\overline{\Phi}_B\colon\frac{L_B\times\R^n}{B}\ra G$.
To see directly that this is a topological isomorphism, define $C'\times\R^n\ra\frac{L_B\times\R^n}{B}$ by $(c,x)\mapsto [([c,0],x)]$.
For $z\in\z^n$ one has $[([\alpha(z),0],z)]=[([0,z],z)]=0$, so this map induces a continuous homomorphism $\Psi_B\colon G\ra\frac{L_B\times\R^n}{B}$.
The maps $\Psi_B$ and $\overline{\Phi}_B$ are inverse to one another.
It follows that $\overline{\Phi}_B$ is a topological isomorphism.
	
Finally, the map $L_B\ra \frac{B}{\z^n}$ by $[c,b]\mapsto b+\z^n$ gives an exact sequence $	0\ra C'\ra L_B\ra \frac{B}{\z^n}\ra 0$.
Since $C'$ is profinite and $\frac{B}{\z^n}$ is a discrete torsion group, $L_B$ is a periodic $\mathbf{LCA}$ group; indeed, for each $\ell\in L_B$ some positive multiple $d\ell$ lies in $C'$, so $\overline{\langle\ell\rangle}$ is contained in finitely many translates of the profinite group $\overline{\langle d\ell\rangle}$.
\end{proof}

Fix $A,B,C,D,E\in \mathbf{LCA}$. 
A continuous homomorphism $f\colon A\ra B$ is {\it proper} if it is open onto its image. 
A {\it proper exact sequence} in $\mathbf{LCA}$ is one for which all maps are proper. 
A short exact sequence $0\ra C\ra D\ra E\ra0$ is a proper short exact sequence in $\mathbf{LCA}$ if and only if $0\ra E^\vee\ra D^\vee\ra C^\vee\ra0$ is a proper short exact sequence in $\mathbf{LCA}$ \cite[Proposition 2.2]{Moskowitz}.
Define a proper short exact sequence $0\ra C\ra D\ra E\ra0$ to be {\it Pontryagin self-dual} if $C\cong E^\vee$, $D\cong D^\vee$, and $E\cong C^\vee$ as topological groups.

\begin{thm}[Pontryagin Self-Dual Proper Short Exact Sequence]\label{self-dual-pses}
Let $G\in\mathbf{AT}$ with $\dima(G)=n$.
There is a periodic group $L\in\mathbf{LCA}$ yielding a Pontryagin self-dual proper short exact sequence in $\mathbf{LCA}${\rm :} $0\ra G^\vee\ra L\times\R^n\ra G\ra0$.
\end{thm}
\begin{proof}
Let $\Delta\in\mathcal{D}(G)$.
By Braconnier's theorem \cite[Theorem 3.3]{HHR} there is a Pontryagin self-dual periodic group $L\in\mathbf{LCA}$ with $\Delta\subseteq L\subseteq \td(G)$.
Applying \cite[Theorem 1.3]{Lewis2019} with	$\widehat{\Delta}_X=L$ and $L(G)=\R^n$ gives a proper short exact sequence $0\ra G^\vee\ra L\times\R^n\ra G\ra 0$.
Its Pontryagin dual $0\ra G^\vee\ra (L\times\R^n)^\vee\ra G\ra 0$ is proper by \cite[Proposition 2.2]{Moskowitz}, where
$(G^\vee)^\vee\cong G$ by Pontryagin reflexivity.
Since $\R^n\cong(\R^n)^\vee$ and $L\cong L^\vee$, one has $(L\times\R^n)^\vee\cong L\times\R^n$.
Hence, the original proper short exact sequence is Pontryagin self-dual.
\end{proof}

Let $H$ be a Hausdorff topological group.
Define the {\it totally disconnected component of $0$ in $H$}, denoted $\td(H)$, to be the set of {\it quasi-torsion elements} $x\in H$: those $x\in H$ such that $\langle x\rangle$ is either finite or its subspace topology is non-discrete and linear.
If $H$ is also complete, as when $H\in\mathbf{LCA}$, then $\td(H)$ coincides with the union, equivalently the sum, of all profinite subgroups of $H$.
If $H$ is compact, then $\td(H)$ is the canonical fully invariant dense divisible totally disconnected component of $0$ in $H$ \cite[Proposition 7.6, Proposition 7.34, Theorem 4.11.6]{DikranjanLewisLothMader}.

Fix $G\in\mathbf{AT}$ with $\dima(G)=n$.
Define the {\it nonarchimedean Lie algebra of $G$}, denoted $\nala(G)$, to be $\Hom(\zhat,G)$ under the compact-open topology.
We prove in Lemma \ref{nala-periodic-lca} below that $\nala(G)$ is a periodic group in $\mathbf{LCA}$.
Lemma \ref{nala-periodic-lca} shows that evaluation at $1$ induces an algebraic isomorphism $\nala(G)\cong\td(G)$; thus, $\nala(G)$ is a divisible periodic $\mathbf{LCA}$ group with $\dimn\nala(G)=n$, uniquely determined up to topological isomorphism by its restricted direct product decomposition \cite[Theorem 3.3]{HHR}.
Define the {\it archimedean Lie algebra of $G$}, denoted $\arla(G)$, to be $\Hom(\R,G)$ under the compact-open topology; then $\arla(G)\cong\R^{\dima(G)}$ as locally compact real topological vector spaces \cite[Theorem 8.22(6)]{HofmannMorris}.
Note that $\dimn\nala(G)=\dima\arla(G)=\dima G=n$.
Set $\az\deff\zhat\times\R$.
Define the {\it adelic Lie algebra of $G$}, denoted $\adla(G)$, to be $\Hom(\az,G)$ under the compact-open topology.

A topological group $G$ {\it has no small subgroups} if there is a neighborhood $U$ of $0$ such that for every subgroup $H\subseteq U$ one has $H=\{0\}$; for example, $\R^n$ and $(\frac{\R}{\z})^n$ have no small subgroups for $n\in\N$.

\begin{lem}\label{nala-periodic-lca}
Let $G\in\mathbf{AT}$ and endow $\Hom(\zhat,G)$ with the compact-open topology.
Then $\Hom(\zhat,G)$ is a periodic group in $\mathbf{LCA}$, and evaluation at $1$ induces an algebraic isomorphism	$\ev_1\colon\Hom(\zhat,G)\ra\td(G)$.
\end{lem}
\begin{proof}
Fix $\Delta\in\mathcal D(G)$ and set $K_\Delta=\{f\in\Hom(\zhat,G)\colon f(\zhat)\subseteq\Delta\}$.
We prove that $K_\Delta$ is a compact open subgroup of $\Hom(\zhat,G)$.
	
First, $K_\Delta$ is open:
Let $\pi_\Delta\colon G\ra\frac{G}{\Delta}$ be the quotient map.
Since $\Delta\in\mathcal D(G)$, one has $\frac{G}{\Delta}\cong(\frac{\R}{\z})^n$, $n=\dima(G)$, and therefore $\frac{G}{\Delta}$ has no small subgroups.
Choose an identity neighborhood $U\subseteq\frac{G}{\Delta}$ containing no nontrivial subgroup.
Then $\{\varphi\in\Hom(\zhat,\frac{G}{\Delta})\colon\varphi(\zhat)\subseteq U\}=\{0\}$.
This is an open neighborhood of $0$ in the compact-open topology, so $\Hom(\zhat,\frac{G}{\Delta})$ is discrete.
Define $\rho_\Delta\colon\Hom(\zhat,G)\ra\Hom(\zhat,\frac{G}{\Delta})$ by $\rho_\Delta(f)=\pi_\Delta\circ f$.
The map $\rho_\Delta$ is continuous and $K_\Delta=\ker(\rho_\Delta)$.
Since the codomain is discrete, $K_\Delta$ is open.
	
We next prove that $K_\Delta$ is compact:
Since $\Delta$ is profinite, its canonical continuous $\zhat$-action shows that evaluation at $1$ induces a topological isomorphism $\ev_1\colon K_\Delta=\Hom(\zhat,\Delta)\ra\Delta$.
Indeed, $\ev_1$ is injective because $\zhat$ is topologically generated by $1$, while its inverse sends $x\in\Delta$ to the continuous homomorphism $z\mapsto z\cdot x$.
Thus, $K_\Delta\cong\Delta$ is compact.
Therefore, $\Hom(\zhat,G)$ has a compact open subgroup and is a locally compact abelian group.
	
It remains to prove periodicity and identify evaluation at $1$.
Let $f\in\Hom(\zhat,G)$.
The homomorphic image $f(\zhat)$ is profinite, so $f(1)\in\td(G)=\bigcup_{\Delta\in\mathcal D(G)}\Delta$ \cite[Theorem 6(1)]{DikranjanLewisLothMader}.
Choose $\Delta_f\in\mathcal D(G)$ such that $f(1)\in\Delta_f$.
Since $f(m)=m f(1)$ for $m\in\Z$ and $\Z$ is dense in $\zhat$, continuity gives $f(z)=z\cdot f(1)\in\Delta_f$ for every $z\in\zhat$.
Hence, $f\in K_{\Delta_f}$.
Thus, every element of $\Hom(\zhat,G)$ belongs to a compact subgroup, so $\Hom(\zhat,G)$ is a periodic group in $\mathbf{LCA}$.

Evaluation at $1$ is injective because $\zhat$ is topologically generated by $1$.
Conversely, if $x\in\td(G)$, then $x$ lies in some $\Delta\in\mathcal D(G)$, and the canonical $\zhat$-action on $\Delta$ defines $f_x\colon\zhat\ra G$ by $f_x(z)=z\cdot x$.
Then $f_x(1)=x$.
Thus, evaluation at $1$ is an algebraic isomorphism onto $\td(G)$.
\end{proof}

Our use of nonarchimedean scalar notation depends fundamentally on \cite[Proposition 4.21]{HHR}: if $L$ is a periodic $\mathbf{LCA}$ group, then the closure of the cyclic subgroup generated by each $g\in L$ is procyclic and has the Sylow decomposition $\overline{\langle g\rangle}\cong\prod_{p\in\p}\overline{\langle g\rangle}_p$.
Consequently, the Sylow decomposition induces a canonical continuous bilinear action $\zhat\times L\ra L$ via $(z,g)\mapsto z\cdot g$, where, for $z=(z_p)_{p\in\p}\in\zhat=\prod_{p\in\p}\z_p$, the $p$-component of $z\cdot g$ is $(z\cdot g)_p=z_pg_p$.
Thus, the scalar notation used below is not an additional multiplication imposed on the additive group $L$, but its canonical external $\zhat$-action.
In particular, delta-subgroups of adelic tori and compact open subgroups of $\nala(G)$ are compact topological $\zhat$-modules.
Moreover, the topological divisible hull $\lowfrown{L}$ of a periodic $\mathbf{LCA}$ group $L$ is unique up to topological isomorphism \cite[Proposition 3.42]{HHR}; for example, $\qhat=\q\otimes\zhat$ is the topological divisible hull of $\zhat$.

\begin{rem}[Archimedean and Nonarchimedean Scalar Actions]
The canonical $\zhat$-action on a periodic $\mathbf{LCA}$ group is the nonarchimedean analogue of the familiar scalar action on the archimedean Lie algebra and its non-locally compact copy ${\rm a}(G)$ in $G$.
In the representation of a finite-dimensional protorus given by the Resolution Theorem \cite[Theorem 8.20]{HofmannMorris}, the archimedean component $\arla(G)\cong\R^n$ carries its natural $\R$-module structure \cite[Theorem 8.22(6)]{HofmannMorris}.
Accordingly, we use $\R$-scalars on archimedean components and $\zhat$-scalars on periodic nonarchimedean components.
On torsion-free divisible periodic $\mathbf{LCA}$ groups, the restricted direct product decomposition extends the canonical $\zhat$-action coordinatewise to a canonical $\qhat$-action.
No additional multiplicative structure is imposed on the ambient $G\in\mathbf{AT}$.
\end{rem}

\begin{thm}[Symmetric Adelic Lie Algebra]\label{symmetric-adelic-lie-algebra}
Let $G\in\mathbf{AT}$ with $\dima(G)=n$.
The adelic Lie algebra of $G$ decomposes algebrotopologically in $\mathbf{LCA}$ as
$\adla(G)\cong\nala(G)\times\arla(G)$.
Here $\nala(G)$ is a divisible periodic $\mathbf{LCA}$ group with $\dimn\nala(G)=n$, uniquely determined up to topological isomorphism by its restricted direct product decomposition \cite[Theorem 3.3]{HHR} in terms of $\q_p$'s, $\Z(p^\infty)$'s, and the $p$-Sylow subgroups of a compact open subgroup, while $\arla(G)\cong\R^n$ is uniquely determined up to topological isomorphism by $n$.
\end{thm}
\begin{proof}
Using the canonical decomposition $\az=\zhat\times\R$, define $\Phi\colon\Hom(\az,G)\ra\Hom(\zhat,G)\times\Hom(\R,G)$ by $\Phi(f)=(f\vert_{\zhat},f\vert_{\R})$.
Its inverse is given by $\Psi(u,v)(z,x)=u(z)+v(x)$.
Thus, $\Phi$ is an algebraic isomorphism.
	
The restriction maps defining $\Phi$ are continuous in the compact-open topology.
To prove continuity of $\Psi$, let $K\subseteq\az$ be compact and let $U$ be an identity neighborhood in $G$.
Choose identity neighborhoods $U_{\rm n},U_{\rm a}\subseteq G$ such that $U_{\rm n}+U_{\rm a}\subseteq U$.
Let $K_{\rm n}\subseteq\zhat$ and $K_{\rm a}\subseteq\R$ be the projections of $K$.
Then $K_{\rm n}$ and $K_{\rm a}$ are compact and
$\Psi([K_{\rm n},U_{\rm n}]\times[K_{\rm a},U_{\rm a}])\subseteq[K,U]$.
Hence, $\Psi$ is continuous, and therefore
$\adla(G)\cong\nala(G)\times\arla(G)$ algebrotopologically.
	
By Lemma \ref{nala-periodic-lca}, $\nala(G)$ is a periodic group in $\mathbf{LCA}$.
Evaluation at $1$ induces an algebraic isomorphism $\nala(G)\cong\td(G)$; hence, $\nala(G)$ is divisible and $\dimn\nala(G)=n$.
The remaining assertions follow from \cite[Theorem 3.3]{HHR} and the topological isomorphism $\arla(G)\cong\R^n$.
\end{proof}

Fix $G\in\mathbf{AT}$ with $\dima(G)=n$.
Define the {\it nonarchimedean exponential} $\exp_{\rm n}\colon\nala(G)\ra G$ by $\exp_{\rm n}(f)=f(1)$.
Define the {\it archimedean exponential} $\exp_{\rm a}\colon\arla(G)\ra G$ by $\exp_{\rm a}(g)=g(1)$.
Define the {\it adelic exponential} $\exp\colon\adla(G)\ra G$ by $\exp(f,g)=f(1)+g(1)$.

The {\it arc component of $0$ in $G\in\mathbf{AT}$}, denoted ${\rm a}(G)$, is the canonical fully invariant dense divisible arc-connected component of $0$ in $G$ \cite[Corollary 8.47]{HofmannMorris} and \cite[Theorem 14, Remark 5]{DikranjanLewisLothMader}.

A rank-$n$ group $B\in\mathbf{DT}$ is {\it strongly indecomposable} if $B$ is not quasi-isomorphic to a direct sum of nontrivial discrete tori of ranks less than $n$; equivalently, $\q\End(B)$ is local \cite[Corollary 7.8]{Arnold}.
Define $G\in\mathbf{AT}$ to be {\it irreducible} if $G^\vee$ is strongly indecomposable; equivalently, $\q\End(G)$ is local \cite[Corollary 7.8]{Arnold}.
We assume that quasi-homomorphism groups and quasi-endomorphism rings of adelic tori consist only of continuous quasi-homomorphisms and are endowed with the compact-open topology.

For an abelian group $A$ and $p\in\p$, one has $\frac{A}{pA}\cong\frac{\z}{p\z}\otimes_\z A$, and we define
$p\text{-rank}(A)\deff\dim_{\z/p\z}\frac{A}{pA}$.
We use the following fact: the delta-subgroups of an irreducible adelic torus $G$ of adelic dimension $(m,n)$ are commensurable, form a semilattice with directed union $\td(G)$, and comprise a set of representatives for the topological isomorphism classes in the isogeny class of a fixed delta-subgroup $C$ satisfying $p\text{-rank}(C)\leq m$ for all but finitely many $p\in\p$ and
$p\text{-rank}(C)\leq n$ for all $p\in\p$.

\begin{thm}[Topological Characterisation via Nonarchimedean Exponential]\label{topological-characterisation-nonarchimedean-exponential}
The nonarchimedean exponential is a continuous algebraic isomorphism of $\nala(G)$ onto $\td(G)$.
If $G\neq\{0\}$, then $\td(G)$ is not locally compact in its subspace topology and $\exp_{\rm n}\colon\nala(G)\ra\td(G)$ is not open.
For each $\Delta\in\mathcal D(G)$, the nonarchimedean exponential maps the compact open subgroup
$K_\Delta=\{f\in\nala(G)\colon f(\zhat)\subseteq\Delta\}$ topologically onto the compact non-open subgroup $\Delta$ of $\td(G)$.
\end{thm}
\begin{proof}
For $f\in\nala(G)$, the homomorphic image $f(\zhat)$ is profinite, and hence $f(1)\in\td(G)$.
Thus, $\exp_{\rm n}$ is well-defined as a map into $\td(G)$.
Evaluation at $1$ is continuous in the compact-open topology.
It is injective because $\zhat$ is topologically generated by $1$.
	
Let $x\in\td(G)$.
Then $x$ lies in a delta-subgroup $\Delta\in\mathcal D(G)$.
The canonical continuous $\zhat$-action on $\Delta$ defines a continuous homomorphism $f_x\colon\zhat\ra G$ by $f_x(z)=z\cdot x$.
Since $\exp_{\rm n}(f_x)=f_x(1)=x$, the map $\exp_{\rm n}$ is surjective onto $\td(G)$.
Thus, $\exp_{\rm n}$ is a continuous algebraic isomorphism of $\nala(G)$ onto $\td(G)$.
	
For $\Delta\in\mathcal D(G)$, Lemma \ref{nala-periodic-lca} shows that $K_\Delta$ is compact and open in $\nala(G)$ and that evaluation at $1$ restricts to a topological isomorphism $K_\Delta\ra\Delta$.
	
Suppose that $\Delta$ is open in the subspace topology on $\td(G)$.
Then there is an identity neighborhood $U$ in $G$ such that	$U\cap\td(G)\subseteq\Delta$.
Since $\td(G)$ is dense in $G$, the set $U\cap\td(G)$ is dense in $U$.
Since $\Delta$ is compact and therefore closed in $G$, it follows that $U\subseteq\Delta$.
Thus, $\Delta$ is open in $G$.
As $G$ is connected, one obtains $\Delta=G$, which is impossible when $G\neq\{0\}$ because $\Delta$ is totally disconnected.
Therefore, $\Delta$ is not open in $\td(G)$, and consequently $\exp_{\rm n}$ is not open.
	
Finally, suppose that $\td(G)$ is locally compact in its subspace topology.
Choose an identity neighborhood $W$ in $\td(G)$ having compact closure $C$ in $\td(G)$.
There is an identity neighborhood $U$ in $G$ such that $U\cap\td(G)\subseteq W\subseteq C$.
The subgroup $\td(G)$ is dense in $G$, while $C$ is compact and therefore closed in $G$.
Hence, $U\subseteq C\subseteq\td(G)$.
Thus, $\td(G)$ is an open subgroup of the connected group $G$, so $\td(G)=G$.
Since $\td(G)$ is totally disconnected, this forces $G=\{0\}$.
Therefore, $\td(G)$ is not locally compact when $G\neq\{0\}$.
\end{proof}

\begin{thm}[Adelic Exponential Yields Lie Category]\label{adelic-exponential-lie-category}
Under the decomposition $\adla(G)\cong \nala(G)\times\arla(G)$, the adelic exponential is the resolution map of \cite[Proposition 6.1, Theorem 6.2, Theorem 6.4]{DikranjanLewisLothMader}, whence continuous, surjective, and open with discrete kernel isomorphic to $\q^n$.
\end{thm}
\begin{proof}
This follows from \cite[Theorem 14]{DikranjanLewisLothMader}.
\end{proof}

The {\it kumu of $G$} is defined to be $\kappa(G)\deff\td(G)\cap{\rm a}(G)$.

\begin{thm}[Kumu the Arbiter]\label{kumu}
The kumu $\kappa(G)$ of $G\in\mathbf{AT}$ is the canonical fully invariant dense subgroup of $G$ algebraically isomorphic to $\q^{\dima G}$.
\end{thm}
\begin{proof}
This follows from \cite[Proposition 6.1, Theorem 6.2, Theorem 6.4]{DikranjanLewisLothMader}.
\end{proof}

Fix $G\in\mathbf{AT}$ with $\dima(G)=n$.
Let $B\subseteq\kappa(G)$ be dense in $G$, and suppose that $A\in\mathbf{DT}$ is isomorphic to $B$.
Fix an isomorphism $\alpha\colon A\ra B$.
The {\it dual topology on $A$ relative to $(G,\alpha)$} is the topology transported to $A$ by $\alpha$, where $B$ carries the subspace topology inherited from $G$.
When no reference to ``relative to $(G,\alpha)$'' is made, we adopt the convention that the dual topology on $A\in\mathbf{DT}$ with no free summands is its dual topology relative to $(A^\vee,\exp_{\rm a}\vert_A)$, under a fixed realization $A\subseteq\q^n\subseteq\R^n$.

For a Hausdorff topological group $K$, its completion is the unique-up-to-topological-isomorphism complete Hausdorff topological group $\widehat K$ admitting an injective continuous homomorphism $i\colon K\ra\widehat K$ that is a topological isomorphism onto the dense subgroup $i(K)$.
One must specify the topology on $K$, since its completion depends upon that topology.
The following is an example illustrating the concept: define the {\it $\z$-adic completion $\widehat E_{\z}$} of a locally free $E\in\mathbf{DT}$ to be the completion of $E$ under the {\it $\z$-adic topology} given by the fundamental system of neighborhoods $\{nE\colon n\in\N\}$ about $0$; then $\widehat E_{\z}$ is well-defined because $E$ locally free $\Rightarrow\bigcap\{nE\colon n\in\N\}=\{0\}$ ($E$ is Hausdorff in the $\z$-adic topology); equivalently, $\widehat E_{\z}\cong\varprojlim_{n\in\N}\frac{E}{nE}$.

\begin{thm}[Dense Subgroups and Dual Topologies]\label{dense-subgroups-dual-topologies}
Let $A\in\mathbf{DT}$ have rank $n\in\N$ and no free summands.
Then the completion of $A$ under its dual topology is topologically isomorphic to $A^\vee$.
\end{thm}
\begin{proof}
Let $H$ be the closure of $A$ in $\R^n$.
By the structure theorem for closed subgroups of $\R^n$, one has
$H=V\oplus\Lambda$, where $V$ is a real vector subspace and $\Lambda$
is discrete and free.
Let $\pi\colon H\ra\Lambda$ be the projection.
Since $A$ is dense in $H$, the subgroup $\pi(A)$ is dense in $\Lambda$.
As $\Lambda$ is discrete, $\pi(A)=\Lambda$.
If $\Lambda\neq\{0\}$, then the exact sequence $0\ra A\cap V\ra A\xrightarrow{\pi}\Lambda\ra 0$ splits because $\Lambda$ is free, giving $A$ a nonzero free summand.
Hence, $\Lambda=\{0\}$.
Since $\q A=\q^n$, the real span of $A$ is $\R^n$, so $V=\R^n$ and $H=\R^n$.
The archimedean exponential $\exp_{\rm a}\colon\R^n\ra A^\vee$ is injective \cite[Corollary 8.47]{HofmannMorris} and satisfies $\exp_{\rm a}(\q^n)=\kappa(A^\vee)$ \cite[Theorem 14]{DikranjanLewisLothMader}.
Set $B=\exp_{\rm a}(A)$.
Then $B\subseteq\kappa(A^\vee)$ and $B\cong A$.
	
Since $A$ is dense in $\R^n$, continuity gives $\exp_{\rm a}(\R^n)\subseteq\overline{B}$.
Thus, ${\rm a}(A^\vee)\subseteq\overline{B}$.
Since ${\rm a}(A^\vee)$ is dense in $A^\vee$, the subgroup $B$ is dense in $A^\vee$.
	
Give $A$ the topology transported from the subspace topology on $B$.
Then $A$ is topologically isomorphic to the dense subgroup $B$ of the compact Hausdorff, and hence complete, group $A^\vee$.
Therefore, $A^\vee$ is the completion of $A$ under its dual topology.
\end{proof}

Let $A\in\mathbf{DT}$ have rank $n$ and no free summands.
Let $F\cong\z^n$ be a full-rank free subgroup of $A$ and set $T\deff\frac{A}{F}$.
Then there is a proper short exact sequence $0\ra F\ra A\ra T\ra0$, whose Pontryagin dual is the proper short exact sequence $0\ra T^\vee\ra A^\vee\ra(\frac{\R}{\z})^n\ra 0$.

Since $\frac{A^\vee}{T^\vee}\cong F^\vee\cong(\frac{\R}{\z})^n$, the profinite group $T^\vee$ is a delta-subgroup of $A^\vee$.
Let $\lowfrown{T^\vee}$ denote its topological divisible hull \cite[Proposition 3.42]{HHR}.
Then $\lowfrown{T^\vee}\times\R^n$ is a minimal divisible locally compact cover of $A^\vee$ \cite[p.13]{Lewis2019}.

In contrast, $\nala(A^\vee)\times\R^n$ is the divisible locally compact cover associated with the adelic exponential, whose diagonal kernel is isomorphic to $\q^n$.
Algebraically, $\nala(A^\vee)\cong\td(A^\vee)$.

Now, fix a number field $K$, that is, a finite extension field of $\q$.
A {\it group of integers of $K$} is a torsion-free subgroup $B_K$ satisfying $\OO_K\subseteq B_K\subseteq K$, $\q B_K=K$, and $\End(B_K)=\OO_K$ \cite[Theorem, p.1]{Zassenhaus}; the terminology is endomorphism-theoretic: the word ``integers'' refers to the endomorphisms of $B_K$, not to the assertion that all elements of $B_K$ are algebraic integers.
Let $\R\cdot K$ denote Zassenhaus' {\it hypercomplex system with the same basis and multiplication constants as $\OO_K$ over the rational number field $\q$}
\cite[Theorem, p.1]{Zassenhaus}.
After choosing a $\z$-basis of $\OO_K$, we identify $\OO_K^\vee$ topologically with $(\R\cdot K)/\OO_K$.
Set $T_K\deff B_K/\OO_K$.

The next theorem is an adelic realisation of Zassenhaus' construction \cite[Theorem, p.1]{Zassenhaus}, not a new proof of that construction. 
The point is that the arithmetic inclusion $\OO_K\subseteq B_K\subseteq K$ becomes, via Pontryagin duality, a distinguished delta-subgroup of an adelic torus together with its archimedean quotient.

\begin{thm}[Adelic Realisation of a Zassenhaus Group of Integers of a Number Field]\label{zassenhaus-realisation}
Every number field $K$, with $n=[K\colon\q]$, contains a $t(\Z)$-homogeneous, strongly indecomposable, locally free torsion-free subgroup $B_K$ such that $\OO_K\subseteq B_K\subseteq K$, $\q B_K=K$, and
$\End(B_K)=\OO_K$.
	
Moreover, $T_K^\vee$ is a delta-subgroup of	$B_K^\vee\in\mathbf{AT}$, and the projection of the diagonal copy of	$\OO_K$ into $T_K^\vee$ is dense. 
After choosing a $\z$-basis of $\OO_K$, there are topological isomorphisms $B_K^\vee\cong	\frac{T_K^\vee\times\R\cdot K}{\OO_K}\cong\frac{\nala(B_K^\vee)\times\R\cdot K}{K}$, and $\frac{B_K^\vee}{T_K^\vee}\cong\frac{\R\cdot K}{\OO_K}\cong(\frac{\R}{\z})^{[K\colon\q]}$.
\end{thm}
\begin{proof}
Apply \cite[Theorem, p.1]{Zassenhaus} with $R=\OO_K$.
This gives a torsion-free subgroup $B_K$ satisfying $\OO_K\subseteq B_K\subseteq K$, $\q B_K=K$, $\End(B_K)=\OO_K$.
The construction gives $B_K$ locally free and $t(\Z)$-homogeneous.
Since $\q\End(B_K)\cong\q\otimes\OO_K\cong K$ is a field, $B_K$ is strongly indecomposable.
Pontryagin duality applied to $0\ra\OO_K\ra B_K\ra T_K\ra 0$ gives the proper short exact sequence
$0\ra T_K^\vee\ra B_K^\vee\ra\OO_K^\vee\ra 0$.
After choosing a $\z$-basis of $\OO_K$, one has $\OO_K^\vee\cong\frac{\R\cdot K}{\OO_K}\cong(\frac{\R}{\z})^n$.
Hence, $T_K^\vee$ is a delta-subgroup of $B_K^\vee$.
By \cite[Theorem 1.3, $Y=\OO_K$]{Lewis2019}, there is a diagonal embedding $\OO_K\hookrightarrow T_K^\vee\times\R\cdot K$ whose projection into $T_K^\vee$ has dense image and for which $B_K^\vee\cong
\frac{T_K^\vee\times\R\cdot K}{\OO_K}$.
The chosen $\z$-basis of $\OO_K$ induces identifications $K\cong\q^n$ and $\R\cdot K\cong\R^n$.
Transporting the adelic exponential sequence along these identifications gives $B_K^\vee\cong\frac{\nala(B_K^\vee)\times\R\cdot K}{K}$ where $K$ denotes the transported diagonal kernel isomorphic to $\q^n$.
\end{proof}

For a choice of group of integers $B_K$ furnished by Theorem \ref{zassenhaus-realisation}, define the
{\it nonarchimedean dimension of the number field $K$} by $\dimn K\deff\dimn(B_K^\vee)$.
The following theorem shows that $\dimn K$, and hence the adelic dimension of $K$, is independent of the choice of $B_K$.

Theorem \ref{zassenhaus-realisation} associates with $K$ an irreducible adelic torus $B_K^\vee$ of archimedean dimension $[K\colon\q]$, whose nonarchimedean structure is encoded by the {\it Richman type} of $B_K$ \cite[p.12, Proposition 1.9]{Arnold}.

\begin{thm}[Nonarchimedean Line through a Number Field]\label{nonarchimedean-line-number-field}
If $K$ is a number field, then $\dimn K=1$.
\end{thm}
\begin{proof}
Let $K/\q$ be a number field with $[K\colon\q]=n$.
Let $B_K$ be a group of integers furnished by Theorem \ref{zassenhaus-realisation}, set $G_K\deff B_K^\vee$, and set $T_K\deff B_K/\OO_K$.
By Theorem \ref{zassenhaus-realisation}, there are resolutions $G_K\cong\frac{\nala(G_K)\times\R\cdot K}{K}$ and $G_K\cong\frac{T_K^\vee\times\R\cdot K}{\OO_K}$.
Choose $b\in \OO_K$ such that $K=\q(b)$. 
Then the primitive element $b$ generates the diagonal copy of $K$ algebrotopologically.
Since that diagonal copy projects densely into the nonarchimedean component of the first resolution, the nonarchimedean line generated by $\q b$ is dense in $\nala(G_K)$, algebraically isomorphic to $\td(G_K)$.
Transport this line through the topological isomorphism relating the two resolutions. 
Its integral part $\z b$ has closure $C\deff\overline{\z b}\subseteq T_K^\vee$ and $C\in\mathcal{D}(G_K)$. 
By construction, $C$ is procyclic. 
Hence, by the definitions, $1=\dimn C=\dimn G_K = \dimn (B_K^\vee) = \dimn K$.
\end{proof}

Theorem \ref{nonarchimedean-line-number-field} says essentially that the (algebraic) primitive element theorem implies the algebrotopological monogenicity of a number field. 

\begin{rem}
The passage from classical monogenicity of number fields to algebrotopological monogenicity is
parallel to the passage from isomorphism to quasi-isomorphism in $\mathbf{DT}$.
Requiring $\OO_K=\mathbb Z[b]$ is an isomorphism-level condition. 
Requiring instead that $\mathbb Z[b]\subseteq\OO_K$	be an order of finite index is the quasi-isomorphism-level condition appropriate to adelic tori.  
After closure in the nonarchimedean component, $\overline{\mathbb Z[b]}$ is a monogenic delta-subgroup commensurable with $\overline{\OO_K}$.
Thus, the obstruction to classical monogenicity is absorbed into the commensurability class of delta-subgroups, just as finite-index discrepancies are absorbed by quasi-isomorphism in $\mathbf{DT}$.
\end{rem}

\begin{example}
We close with a Pontryagin-duality computation illustrating that the dual of a familiar locally compact group need not lie wholly on either side of the usual discrete–compact correspondence.
We show $(\A/\z)^\vee \cong \zhat + \q$, the sum of a compact group and a discrete torsion-free group.
We have $\frac{\A}{\zhat+\q}\cong \frac{(\A/\zhat)}{\q}\cong\frac{(\q/\z)\times\R}{\q}\cong\frac{\R}{\z}$ so there is a proper short exact sequence $0\ra\zhat+\q\ra\A\ra\frac{\R}{\z}\ra 0$ in $\mathbf{LCA}$ with Pontryagin dual proper short exact sequence $0\ra \z\ra\A\ra(\zhat+\q)^\vee\ra 0$ in $\mathbf{LCA}$, so $\frac{\A}{\z}\cong (\zhat+\q)^\vee$; equivalently, $(\frac{\A}{\z})^\vee\cong \zhat+\q$.
\end{example}

\end{document}